# An Iterative Task-Driven Framework for Resilient LiDAR Place Recognition in Adverse Weather

### v1

@@ III-C LiDAR Place Recognition module
$\hat{\mathbf{F}}_{\text{s}}=\text{Concat}(\mathbf{A}_{\text{s}}^{1},\mathbf{A}_{\text{s}}^{2},...,\mathbf{A}_{\text{s}}^{m})$ (20)
@@
$\mathbf{F}{{}_{\text{m}}}=\text{IDWT}(\hat{\mathbf{F}}_{\text{LL}},\hat{\mathbf{F}}_{\text{LH}},\hat{\mathbf{F}}_{\text{HL}},\hat{\mathbf{F}}_{\text{HH}})+\mathbf{F}_{\text{0}}$ (21)
@@
8:              $\theta_{LPR}\leftarrow Adam(\theta_{LPR},g_{LPR},\ell_{LPR})$
@@
13:              $\theta_{LDR}\leftarrow Adam(\theta_{LDR},g_{LDR},\ell_{LDR})$
@@ IV Experiments
Table II: Benchmarking results of LPR methods on Weather-KITTI. Top: Direct evaluation on degraded data. Middle: Separate training with ITDNet-D. Bottom: Joint iterative training of the full ITDNet. ITDNet-D and ITDNet-P denote the LDR and LPR modules of ITDNet, respectively.
Methods Seq 00 Seq 07
@@ IV-A Dataset
Boreas [38] provides over 350 km of driving data collected along a repeated route over one year, covering real-world seasonal and weather variations, including various levels of rain and snow. The platform is equipped with a 128-beam LiDAR and GPS/IMU sensors that offer high-precision global ground-truth poses. We select five representative sequences from Boreas; all sequences are recorded along a similar route under different seasons and weather conditions, with their weather conditions summarized in Table I. For the LDR tasks, clean ground-truth scans are unavailable for degraded scans captured in adverse weather. To address this, we generate training pairs by projecting each degraded scan from Seq 02-03 onto a spatially aligned clean scan from Seq 00. Specifically, for a scan $\text{S}_{2}$ with pose $(\mathbf{R}_{2},\mathbf{T}_{2})$ in Seq 02, we identify the closest corresponding scan $\text{S}_{0}$ with pose $(\mathbf{R}_{0},\mathbf{T}_{0})$ from Seq 00, under the constraint that the spatial distance between the two poses is less than 0.01m and the angular difference is below $0.1^{\circ}$. Each point $\mathbf{p}_{2}$ in $\text{S}_{2}$ is then projected into the frame of $\text{S}_{0}$ via:
$\mathbf{p}_{0}=\mathbf{R}_{0}^{-1}\left(\mathbf{R}_{2}\mathbf{p}_{2}+\mathbf{T}_{2}-\mathbf{T}_{0}\right)$ (26)
This process generates clean-degraded pairs for LDR training. While environmental changes may lead to some misalignment, we emphasize that our goal is not to recover exact point-to-point fidelity, but rather to enhance feature consistency for downstream LPR tasks. Therefore, this approximation is sufficient for training, as demonstrated by the improved place recognition performance in the subsequent LPR testing results. For the LPR task, we train on Seq 00, use Seq 01 as the test database, and Seq 04, which contains the most severe weather degradations, as the test query set. Since each sequence includes both forward and reverse traversals, we define two evaluation settings to assess long-term place recognition performance: an easy mode using only the forward traversal of the query set, and a hard mode incorporating both directions to introduce greater viewpoint variation and challenge.
@@ IV-B Experimental Setup
In our experiments, we first project the point clouds onto LiDAR range images. The input range images $\mathcal{I}_{0}$ are sized $64\times 1440\times 2$ for Weather-KITTI [11] and Weather-Apollo, and $128\times 1920\times 2$ for Boreas [38], reflecting the number of LiDAR beams and the average point count per scan in each dataset. Both the LDR and LPR modules are trained for 200 epochs using the AdamW optimizer [42] with an initial learning rate of $10^{-4}$, decayed to $10^{-6}$ via cosine annealing ($t_{\text{max}}=100$). Training samples are cropped to $32\times 480$ patches with random flipping for data augmentation. The loss balance weight $\lambda$ for the LDR task is set to 0.1. All experiments are run on an Ubuntu 18.04 system with dual Intel Xeon 8280 CPUs, four NVIDIA V100 GPUs, and 32 GB RAM.
@@
Method Snow Fog Rain
@@
ITDNet 0.66 0.86 0.97 0.81 0.67 0.85 0.96 0.80 0.69 0.87 0.97 0.82
Table IV: Benchmarking results of LPR methods on Real World Dataset Boreas.
Method Seq 04 (Easy) Seq 04 (Hard)
@@ IV-C Evaluation for the LPR task
In this section, we compare our proposed ITDNet with three state-of-the-art LPR methods: the image-based OverlapTransformer (OT) [23] and CVTNet[1], and the point-based LCDNet[37]. Evaluations are conducted under adverse weather conditions in the following three setting: 1) Direct: each LPR model, including our our LPR module (ITDNet-P), is evaluated directly on degraded dataset; 2) Separate: each LPR model is combined with our LDR module (ITDNet-D) via separate training; 3) Union: our full ITDNet pipeline is trained end-to-end through joint iterative optimization of LDR and LPR modules.
@@ V Conclusion and Future Work
This paper proposes ITDNet, an end-to-end framework that integrates LiDAR data restoration (LDR) and place recognition (LPR) through an iterative learning strategy to address LiDAR degradation under adverse weather. The LDR module employs Dual-Domain Mixer (DDM) and Semantic-Aware Generator (SAG) to recover structural and semantic features, while the LPR module utilizes Multi-scale Frequency Transformer (MFT) and Wavelet Pyramid NetVLAD (WPN) for robust global descriptor extraction. Extensive experiments on Weather-KITTI, Weather-Apollo, and Boreas demonstrate that ITDNet significantly enhances both restoration quality and recognition performance. We also introduce the Feature Similarity Score (FSS), which strongly correlates with downstream LPR accuracy. Ablation studies further verify the effectiveness of each component. In future work, we aim to explore unsupervised solutions for LiDAR perception under adverse weather, addressing the challenge of limited supervision in real-world scenarios.

### v2

@@ III-C LiDAR Place Recognition module
$\hat{\mathbf{F}}_{\text{s}}=\text{Concat}(\mathbf{A}_{\text{s}}^{1},\mathbf{A}_{\text{s}}^{2},...,\mathbf{A}_{\text{s}}^{m}).$ (20)
@@
$\mathbf{F}{{}_{\text{m}}}=\text{IDWT}(\hat{\mathbf{F}}_{\text{LL}},\hat{\mathbf{F}}_{\text{LH}},\hat{\mathbf{F}}_{\text{HL}},\hat{\mathbf{F}}_{\text{HH}})+\mathbf{F}_{\text{0}}.$ (21)
@@
8:              $\theta_{LPR}\leftarrow AdamW(\theta_{LPR},g_{LPR},\ell_{LPR})$
@@
13:              $\theta_{LDR}\leftarrow AdamW(\theta_{LDR},g_{LDR},\ell_{LDR})$
@@ IV Experiments
Table II: Benchmarking results of LPR methods on Weather-KITTI. ITDNet-D and ITDNet-P denote the LDR and LPR modules of ITDNet, respectively.
Strategy Methods Seq 00 Seq 07
@@ IV-A Dataset
Boreas [46] provides over 350 km of driving data collected along a repeated route over one year, covering real-world seasonal and weather variations, including various levels of rain and snow. The platform is equipped with a 128-beam LiDAR and GPS/IMU sensors that offer high-precision ground-truth poses. We select five representative sequences from Boreas; all sequences are recorded along a similar route under different seasons and weather conditions, with their weather conditions summarized in Table I. For the LDR tasks, clean ground-truth scans are unavailable for degraded scans captured in adverse weather. To address this, we generate training pairs by projecting each degraded scan from Seq 02-03 onto a spatially aligned clean scan from Seq 00. Specifically, for a scan $\text{S}_{2}$ with pose $(\mathbf{R}_{2},\mathbf{T}_{2})$ in Seq 02, we identify the closest corresponding scan $\text{S}_{0}$ with pose $(\mathbf{R}_{0},\mathbf{T}_{0})$ from Seq 00, under the constraint that the spatial distance between the two poses is less than 0.01m and the angular difference is below $0.1^{\circ}$. Each point $\mathbf{p}_{2}$ in $\text{S}_{2}$ is then projected into the frame of $\text{S}_{0}$ via:
$\mathbf{p}_{0}=\mathbf{R}_{0}^{-1}\left(\mathbf{R}_{2}\mathbf{p}_{2}+\mathbf{T}_{2}-\mathbf{T}_{0}\right).$ (26)
This process generates clean-degraded pairs for LDR training. While environmental changes may lead to some misalignment, we emphasize that our goal is not to recover exact point-to-point fidelity, but rather to enhance feature consistency for downstream LPR tasks. Therefore, this approximation proves sufficient for training, as evidenced by the improved place recognition performance in Table LABEL:boreas_results. For the LPR task, we train on Seq 00, use Seq 01 as the test database, and Seq 04, which contains the most severe weather degradations, as the test query set. Since each sequence includes both forward and reverse traversals, we define two evaluation settings to assess long-term place recognition performance: an easy mode using only the forward traversal of the query set, and a hard mode incorporating both directions to introduce greater viewpoint variation and challenge.
@@ IV-B Experimental Setup
In our experiments, we first project the point clouds onto LiDAR range images. The input range images $\mathcal{I}_{0}$ are sized $64\times 1440\times 2$ for Weather-KITTI [13] and Weather-Apollo, and $128\times 1920\times 2$ for Boreas [46], reflecting the number of LiDAR beams and the average point count per scan in each dataset. Both the LDR and LPR modules are trained for 200 epochs using the AdamW optimizer [50] with an initial learning rate of $10^{-4}$, decayed to $10^{-6}$ via cosine annealing ($t_{\text{max}}=100$). Training samples are cropped to $32\times 480$ patches with random flipping for data augmentation. To mitigate the impact of noisy pseudo-labels generated by the early-stage LPR model, the loss weight $\lambda$ for the LDR task is set to 0.01 for the first 30 epochs. After this warm-up period, $\lambda$ is increased to 0.1 to strengthen task-driven supervision. All experiments are run on an Ubuntu 18.04 system with dual Intel Xeon 8280 CPUs, four NVIDIA V100 GPUs, and 32 GB RAM.
@@
Strategy Method Snow Fog Rain
@@
Union ITDNet 0.66 0.86 0.97 0.81 0.67 0.85 0.96 0.80 0.69 0.87 0.97 0.82
Table IV: Benchmarking results of LPR methods on Boreas.
Strategy Method Seq 04 (Easy) Seq 04 (Hard)
@@ IV-C Evaluation for the LPR task
In this section, we compare our proposed ITDNet with three state-of-the-art LPR methods: the image-based OverlapTransformer (OT) [26] and CVTNet[1], and the point-based LCDNet[45]. Evaluations are conducted under adverse weather conditions in the following three setting: 1) Direct: each LPR model, including our LPR module (ITDNet-P), is evaluated directly on degraded dataset; 2) Separate: each LPR model is combined with our LDR module (ITDNet-D) via separate training; 3) Union: our full ITDNet pipeline is trained end-to-end through joint iterative optimization of LDR and LPR modules.
@@ V Conclusion and Future Work
This paper presents ITDNet, an end-to-end framework that jointly performs LiDAR data restoration (LDR) and place recognition (LPR) through an iterative learning strategy to address LiDAR degradation in adverse weather. The LDR module leverages a Dual-Domain Mixer (DDM) and Semantic-Aware Generator (SAG) to recover structural and semantic features, while the LPR module employs a Multi-scale Frequency Transformer (MFT) and Wavelet Pyramid NetVLAD (WPN) to extract robust global descriptors. Experiments on Weather-KITTI, Weather-Apollo, and Boreas demonstrate clear improvements in both restoration quality and recognition accuracy. We also introduce the Feature Similarity Score (FSS), which correlates well with downstream LPR performance. Ablation studies confirm the effectiveness of each component. Despite these advances, limited supervision and the sim-to-real gap remain key challenges. Future work will explore unsupervised learning and more generalizable models for LiDAR perception under adverse weather.